# Rehabilitation Exercise Quality Assessment through Supervised Contrastive Learning with Hard and Soft Negatives

### v1

@@ 4.4 Experimental Results
UI-PRMD encoder only 0.9450 0.9100 0.8350 0.8100 0.9350 0.8400 0.9000 0.8850 0.8950 0.9750
encoder + projection head 0.9900 0.9650 0.9750 0.9850 0.9800 0.9750 0.9800 0.9850 0.9900 0.9950
@@
4.5 Transfer Learning to KIMORE
The IRDS [26] and KIMORE [18] datasets, both acquired using the Kinect One sensor, feature structurally similar body-joint data; specifically, the adjacency matrices internal to the body joint graph are identical. Therefore, an ST-GCN encoder pre-trained on IRDS [26] is expected to maintain spatial-temporal relationships when transferred to KIMORE [18]. This encoder, alongside the projection head, undergoes training using the proposed method on IRDS [26], as outlined in subsection 4.3. The pre-trained ST-GCN encoder is kept, a feed-forward regression layer is added to it, and then fine-tuned to individual exercise types in KIMORE [18]. Alternatively, an untrained ST-GCN is created, a feed-forward regression layer is added to it, and trained from scratch. Table 5 presents Spearman’s rank correlations for individual exercise types in KIMORE [18] and compares them with Spearman’s rank correlations for four previous methods. As indicated in Table 5, utilizing transfer learning improves outcomes over training a model from scratch. Except for k02, the approach employing transfer learning outperforms earlier methods for the remaining exercises. Fig. 4 illustrates the training and validation Mean Squared Error (MSE) loss curves of the proposed method across consecutive epochs for the five exercise types in the KIMORE dataset [18]. While Fig. 4 (a), (b), and (d) demonstrate that the pre-trained ST-GCN achieves improved results, this advantage does not extend to Fig. 4 (c) and (e), where the effectiveness of transfer learning is reduced due to exercise misalignment. Certain exercise types in KIMORE [18] do not closely align with those in IRDS [26]. For example, there are no IRDS movements/exercises that encompass squatting, Fig. 4 (e). Moreover, IRDS mandates that the torso remain stationary in all exercises, which excludes any trunk rotations, Fig. 4 (c). However, the lifting of arms, as shown in Fig. 4 (a), may include IRDS movements i03 to i07. Generally, the pre-trained ST-GCN outperforms the vanilla ST-GCN in exercises that share movements with the source dataset. Where there is little or no overlap between the target and source, the performance of the pre-trained ST-GCN should be, at the very least, comparable to that of the un-trained ST-GCN.
@@ 5 Conclusion and Future Works
Our research led to the development of a novel supervised contrastive learning framework for rehabilitation exercise quality assessment. This framework adeptly utilizes entire datasets to train a single, versatile model, effectively addressing the challenge of limited samples for individual exercise types in rehabilitation exercise datasets [25, 26, 18]. The successful application of the proposed framework to three publicly available rehabilitation exercise datasets confirms its efficacy and establishes a new standard in accuracy, outperforming existing methods. The proposed model’s increased generalizability and reduced parameter count are notable advancements, enhancing efficiency and streamlining integration into practical exercise-based virtual rehabilitation platforms. However, there are some limitations in our work. Similar to previous approaches in this field, it requires a preliminary exercise-type classification model before conducting an exercise quality assessment. A significant advancement to the proposed method would be the development of a multitask model capable of simultaneous exercise type classification and assessment, further simplifying the assessment process. Future research directions should include diversifying dataset types and sizes, applying this framework across more varied rehabilitation scenarios, and refining the model for a greater variety of exercise types. Another major area for improvement is adding interpretability to our method, through techniques such as Gradient-based Class Activation Maps [23] for ST-GCNs [54]. This would facilitate an understanding of which body joints and specific timestamps contribute to incorrect exercises. This could then be translated into visual or textual feedback for patients [55], enhancing the utility and effectiveness of virtual rehabilitation programs.

### v2

@@ 4.4.2 Impact of Retaining the Projection Head
UI-PRMD Proposed (encoder only) 0.9450 0.9050 0.8350 0.8400 0.9450 0.8450 0.9000 0.8850 0.8950 0.9750
Proposed (encoder + projection head) 0.9900 0.9850 0.9850 0.9850 0.9800 0.9750 0.9800 0.9850 0.9900 0.9950
ST-GCN with RI [23] 0.9900 0.9900 0.9750 0.9750 0.9800 0.9900 0.9900 0.9800 0.9900 0.9950
Proposed (encoder + projection head)
and SVM as binary classifier
1.0000 1.0000 0.9550 0.9800 0.9450 0.9700 0.9800 0.9700 0.9700 0.9750
4.4.3 Visualization of Learned Representations
@@
Figure 3: t-SNE visualization of representations learned through the proposed supervised contrastive learning approach for (a) UI-PRMD and (b) IRDS datasets. Representations are color-coded on an exercise basis, and ”+” denotes the reference representations, i.e., cluster centers.
4.4.4 Number of Parameters Compared to Previous Works
@@
4.4.5 Transfer Learning to KIMORE
The IRDS [26] and KIMORE [18] datasets, both acquired using the Kinect One sensor, feature structurally similar body-joint data; specifically, the adjacency matrices internal to the body joint graph are identical. Therefore, an ST-GCN encoder pre-trained on IRDS [26] is expected to maintain spatial-temporal relationships when transferred to KIMORE [18]. This encoder, alongside the projection head, undergoes training using the proposed method on IRDS [26], as outlined in subsection 4.3.
As explained in subsection 4.2, the exercise quality assessment problem in IRDS [26] is a binary classification, categorizing exercises as either correct or incorrect. In contrast, in KIMORE [18], the problem is treated as a regression task, inferring a real-valued number ranging from 0 to 50. To address this inconsistency, while the ST-GCN encoder pre-trained on IRDS [26] is retained, the projection head pre-trained on IRDS [26] is replaced with a two-layer fully connected regression network, transforming embeddings from a dimensionality of 256 to 128 and then to a real-valued number. The fully connected regression layer is fine-tuned to individual exercise types in KIMORE [18]. In another setting, an untrained ST-GCN encoder is created, a two-layer fully connected regression network with the same structure as the previous setting is added to it, and trained from scratch on individual exercise types in KIMORE [18].
Table 5 presents Spearman’s rank correlations for individual exercise types in KIMORE [18] and compares them with Spearman’s rank correlations for four previous methods. The correlations are calculated between the models’ real-valued outputs and the ground-truth exercise quality annotations in the dataset. As shown in Table 5, utilizing transfer learning improves outcomes over training a model from scratch. Except for k02, the approach employing transfer learning outperforms earlier methods for the remaining exercises.
Fig. 4 illustrates the training and validation Mean Squared Error (MSE) loss curves of the proposed method across consecutive epochs for the five exercise types in KIMORE [18]. While Fig. 4 (a), (b), and (d) demonstrate that the pre-trained ST-GCN achieves improved results, this advantage does not extend to Fig. 4 (c) and (e), where the effectiveness of transfer learning is reduced due to exercise misalignment. Certain exercise types in KIMORE [18] do not closely align with those in IRDS [26]. For example, there are no IRDS movements/exercises that encompass squatting, Fig. 4 (e). Moreover, IRDS mandates that the torso remain stationary in all exercises, which excludes any trunk rotations, Fig. 4 (c). However, the lifting of arms, as shown in Fig. 4 (a), may include IRDS movements i03 to i07. Generally, the pre-trained ST-GCN outperforms the un-trained ST-GCN in exercises that share movements with the source dataset. Where there is little or no overlap between the target and source, the performance of the pre-trained ST-GCN should be, at the very least, comparable to that of the un-trained ST-GCN.
@@ 4.4.5 Transfer Learning to KIMORE
Figure 4: Training and validation Mean Squared Error (MSE) loss curves of the proposed method across consecutive epochs for the five exercise types in the KIMORE dataset: (a) lifting of arms, (b) trunk lateral tilt, (c) trunk rotation, (d) pelvis rotation, and (e) squatting. The training was performed in two different settings: training an untrained ST-GCN encoder from scratch, and fine-tuning an ST-GCN encoder pre-trained on the IRDS dataset.
@@ 5 Conclusion and Future Works
Our research led to the development of a novel supervised contrastive learning framework for rehabilitation exercise quality assessment. This framework effectively utilizes entire datasets to train a single, versatile model, effectively addressing the challenge of limited samples for individual exercise types in rehabilitation exercise datasets. The successful application of the proposed framework to three publicly available rehabilitation exercise datasets confirms its efficacy and establishes a new standard in accuracy, outperforming existing methods. The proposed model’s increased generalizability and reduced parameter count are notable advancements, enhancing efficiency and streamlining integration into practical exercise-based virtual rehabilitation platforms. However, there are some limitations in our work. Similar to previous approaches in this field, it requires a preliminary exercise type classification model before conducting an exercise quality assessment. A significant advancement in the proposed method would be the development of a multitask model capable of simultaneous exercise type classification and assessment, further simplifying the assessment process. Future research directions should include applying this framework across more varied rehabilitation scenarios and refining the model for a greater variety of exercise types. Another major area for improvement is adding interpretability to our method, through techniques such as gradient-based class activation maps for ST-GCNs. This would facilitate an understanding of which body joints and specific timestamps contribute to incorrect exercises. This could then be translated into visual or textual feedback for patients, enhancing the utility and effectiveness of virtual rehabilitation programs.
@@
Funding
This research was funded by New Frontiers in Research Fund, Canada and TRANSFORM HF Undergraduate Summer Research Program, Canada.
Author Biography
Mark Karlov:
B.Sc. in Electrical and Computer Engineering from University of Toronto. Expert in deep learning for data analysis across various modalities.
Ali Abedi: Ph.D. in Electrical Engineering and Computer Science, Postdoctoral Fellow at U of T and UHN. Specializes in machine learning and deep learning technologies.
Shehroz S. Khan: Ph.D. in Computer Science, Scientist at UHN, Assistant Professor at U of T. Focuses on AI applications in healthcare and rehabilitation.